It costs £11.25 a day to stay in a hotel in
September and £88.25 a week in August.by
comparing the cost of a week holiday in
September with that of a week's holiday in
August find ratio of cost in lowest terms

Answer by greenestamps(13200)



Since the problem asks to compare the weekly costs as a ratio in lowest terms, it might be easiest to convert the given decimal costs into fractions. Note this will be relatively easy, because both given rates involve .25 = 1/4 of a pound.

one day in September: 11.25 = 45/4
one week = 7 days in September: 7*(45/4) = 315/4
one week in August: 88.25 = 353/4

The ratio is 315/353. Those whole numbers have no common factor, so that ratio is in lowest terms.

ANSWER: 315/353, or 315:353